40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 29 17:10:09 2021, Anchal, Summary, CDS, c1teststand problems summary c1teststand_issues_summary.pdf
    Reply  Thu Sep 30 14:09:37 2021, Anchal, Summary, CDS, New way to ssh into c1teststand 
       Reply  Thu Mar 3 15:37:40 2022, Anchal, Summary, CDS, c1teststand restructured 
    Reply  Mon Oct 4 11:05:44 2021, Anchal, Summary, CDS, c1teststand problems summary 
       Reply  Mon Oct 4 18:00:16 2021, Koji, Summary, CDS, c1teststand problems summary 
          Reply  Tue Oct 5 17:58:52 2021, Anchal, Summary, CDS, c1teststand problems summary 
       Reply  Tue Oct 5 18:00:53 2021, Anchal, Summary, CDS, c1teststand time synchronization working now 
          Reply  Mon Oct 11 17:31:25 2021, Anchal, Summary, CDS, Fixed mounting of mx devices in fb. daqd_dc is running now. 
             Reply  Mon Oct 11 18:29:35 2021, Anchal, Summary, CDS, Moving forward? 
                Reply  Tue Oct 12 17:20:12 2021, Anchal, Summary, CDS, Connected c1sus2 to martian network 
                   Reply  Tue Oct 12 23:42:56 2021, Koji, Summary, CDS, Connected c1sus2 to martian network 
                      Reply  Wed Oct 13 11:25:14 2021, Anchal, Summary, CDS, Ran c1sus2 models in martian CDS. All good! CDS_screens_running.png
                         Reply  Tue Oct 19 18:20:33 2021, Ian MacMillan, Summary, CDS, c1sus2 DAC to ADC test data3_Plots.pdfdata2_Plots.pdf
                            Reply  Tue Oct 19 23:43:09 2021, Koji, Summary, CDS, c1sus2 DAC to ADC test P_20211019_224433.jpgP_20211019_224122.jpgP_20211019_224400.jpgP_20211019_224411.jpg
                               Reply  Wed Oct 20 11:48:27 2021, Anchal, Summary, CDS, Power supple configured correctly. 
                               Reply  Tue Oct 26 18:24:00 2021, Ian MacMillan, Summary, CDS, c1sus2 DAC to ADC test data2_Plots.pdfdata3_Plots.pdf
                         Reply  Wed Dec 22 17:40:22 2021, Anchal, Summary, CDS, c1su2 model updated with SUS damping blocks for 7 SOSs 
                            Reply  Wed Dec 29 20:09:40 2021, rana, Summary, CDS, c1su2 model updated with SUS damping blocks for 7 SOSs 
                            Reply  Fri Mar 4 11:04:34 2022, Anchal, Summary, CDS, c1susaux2 system setup and running 40mBHD_C1SUSAUX2_Acromag_Chassis.pdf
                               Reply  Mon Mar 7 19:38:47 2022, Anchal, Summary, CDS, c1susaux2 slow controls issues 
                               Reply  Mon Mar 14 12:20:05 2022, Anchal, Summary, CDS, c1susaux2 slow controls acromag chassis installed BHD_WatchDogs.png40mBHD_C1SUSAUX2_Acromag_Chassis.pdf
                                  Reply  Thu Mar 17 19:12:44 2022, Anchal, Summary, CDS, c1auxey1 slow controls acromag chassis installed, not powered 
                                     Reply  Fri Mar 18 18:39:13 2022, Yehonathan, Summary, CDS, c1auxey1 slow controls acromag chassis installed, powered 
                                        Reply  Fri Mar 18 19:10:51 2022, Anchal, Summary, CDS, c1auxey1 slow controls issues 
                                           Reply  Mon Mar 21 18:42:06 2022, Anchal, Summary, CDS, c1auxey1 slow controls issues 
                                  Reply  Mon Apr 4 17:03:47 2022, Anchal, Summary, CDS, c1susaux2 slow controls acromag chassis fixed and installed Screenshot_2022-04-04_17-03-26.png
                            Reply  Tue Mar 15 11:52:34 2022, Anchal, Summary, CDS, c1su2 model updated for sending Run/Acquire Binary Output to Binary Interface card c1su2.pdf
                               Reply  Tue Mar 15 14:10:41 2022, Anchal, Summary, CDS, c1su2 model remade, reinstalled, restarted after the update 
             Reply  Tue Oct 12 17:10:56 2021, Anchal, Summary, CDS, Some more information 
Message ID: 16415     Entry time: Tue Oct 19 23:43:09 2021     In reply to: 16414     Reply to this: 16417   16430
Author: Koji 
Type: Summary 
Category: CDS 
Subject: c1sus2 DAC to ADC test 

(Because of a totally unrelated reason) I was checking the electronics units for the upgrade. And I realized that the electronics units at the test stand have not been properly powered.

I found that the AA/AI stack at the test stand (Attachment 1) has an unusual powering configuration (Attachment 2).
- Only the positive power supply was used / - The supply voltage is only +15V / - The GND reference is not connected to anywhere.

For confirmation, I checked the voltage across the DC power strip (Attachments 3/4). The positive was +5.3V and the negative was -9.4V. This is subject to change depending on the earth potential.

This is not a good condition at all. The asymmetric powering of the circuit may cause damages to the opamps. So I turned off the switches of the units.

The power configuration should be immediately corrected.

  1. Use both positive and negative supply (2 power supply channels) to produce the positive and the negative voltage potentials. Connect the reference potential to the earth post of the power supply.
    https://www.youtube.com/watch?v=9_6ecyf6K40   [Dual Power Supply Connection / Serial plus minus electronics laboratory PS with center tap]
  2. These units have DC power regulator which produces +/-15V out of +/-18V. So the DC power supplies are supposed to be set at +18V.

 

Attachment 1: P_20211019_224433.jpg  3.889 MB  Uploaded Wed Oct 20 00:46:30 2021  | Hide | Hide all | Show all
P_20211019_224433.jpg
Attachment 2: P_20211019_224122.jpg  3.843 MB  Uploaded Wed Oct 20 00:49:22 2021  | Hide | Hide all | Show all
P_20211019_224122.jpg
Attachment 3: P_20211019_224400.jpg  4.071 MB  Uploaded Wed Oct 20 00:49:27 2021  | Show | Hide all | Show all
Attachment 4: P_20211019_224411.jpg  4.405 MB  Uploaded Wed Oct 20 00:49:27 2021  | Show | Hide all | Show all
ELOG V3.1.3-